OpenStudy (anonymous):

Find the functions which satisfy the following conditions:

OpenStudy (anonymous):

\[f(x+y) + f(x-y) = 2f(x)f(y) \] And that: \[\lim_{x \rightarrow \infty} f(x) = 0 \]

OpenStudy (anonymous):

f : R->R ???

OpenStudy (anonymous):

It is cos(x)

OpenStudy (anonymous):

@mukushla Yes.

OpenStudy (anonymous):

And x,y are real.

OpenStudy (anonymous):

let x=y=0\[2f(0)=2f^2(0)\]so f(0)=0 or 1

OpenStudy (anonymous):

f(x) = 0 Would get us a constant function. That is one solution.

OpenStudy (anonymous):

yeah let suppose f(0)=1

OpenStudy (anonymous):

What about f(x) = cos(x)?

OpenStudy (anonymous):

No. Because, \[\lim_{x \rightarrow \infty} cosx \] is not 0.

OpenStudy (anonymous):

It is zero

OpenStudy (anonymous):

@mukushla How did you get that from x = 0? @Zekarias No, it can be anything from zero to one.

OpenStudy (anonymous):

oh my bad sorry :\

OpenStudy (anonymous):

** @Zekarias -1 to 1 sorry.

OpenStudy (anonymous):

setting x=y it gives\[f(2x)+1=2f^2(x)\]and y=-x gives\[f(2x)+1=2f(x)f(-x)\]so \[f(x)=f(-x)\]

OpenStudy (anonymous):

is this correct ? and how it helps ... i dont know yet

OpenStudy (experimentx):

how about we use the same technique as we used yesterday.

OpenStudy (anonymous):

i think the only solution is f(x)=0

OpenStudy (experimentx):

that's trivial ... let's hunt for some.

OpenStudy (experimentx):